OpenStudy (anonymous):

2x+4y=5 4x+2y=8 Determine whether the pair of lines are perpendicular, parallel, or neither...Thanks!

OpenStudy (anonymous):

neither

OpenStudy (anonymous):

slope of 1st line is -2/4=-1/2 slope of 2nd line is -4/2=-2 as slope of both line is not equal hence they are not parellel if product of slope is -1 then both lines are perpendicular that is not the case hence neither